OpenStudy (anonymous):

]Identify the correct slope and y intercept of the equation 6x - 3y = 12. slope = 2; y intercept at (0, -4) slope = -2; y intercept at (0, 4) slope = 1/2; y intercept at (0, 15) slope =1/2 ; y intercept at (0, 9)

OpenStudy (anonymous):

solve the equation for y @christianfilms11

OpenStudy (anonymous):

do you know how to?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

6x - 3y = 12 first i would say to make it easier you can move the -3y to the other side, because negatives will get confusing so moving the -3y over you get 6x=12+3y then you want to move the +12 to the left 6x-12=3y now you divide both 6x and -12 by 3 to get rid of the 3 from the 3y, which will leave you with just y \[\frac{ 6x }{ 3 }-\frac{ 12 }{ 3 }=y\] then simplify 2x-4=y

OpenStudy (anonymous):

sooooo from the point slope formula -> y=mx+b mx is the slope and b is the y intercept

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

good job!

OpenStudy (anonymous):

learn how to solve for y or x in any equation. this will come in handy in higher level math classes

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

no problem